A daughter is 28 years younger than her father. in five years, the father will be 3 times as old as his daughter. How old is each now?
Answer by usyim88hk(158)

1)let the age of the daughter be x
2)x + 28 is the age of her dad
3)so x+5 will be the age of the daughter after 5 years
4)and 3(x+5) will be the age of the dad after three years which is also equal to 3x+15
5)x+33 is also the age of her dad after five years
6)so x+33 = 3x+15 (both the age of her father after 5 years)
7)x+33-15=3x+15-15
8)x+18=3x
9)x+18-x=3x-x
10)18=2x
11)x=9(So the age of the girl now is 9 years old)
12)the age of her father is x+28 so he is 37 years old now
------------------------------------------------------------------------------
Answer: age of the daughter is 9 and age of the dad is 37